# On heavy-flavour jets with Soft Drop

### 18 Dec 2023 v1

@@ Appendix A Running coupling integrals with quark-mass thresholds
Throughout the paper, we have made use of the strong coupling in the Catani-Marchesini-Webber (CMW) scheme: Catani:1990rr
@@
The limits of integration, as well as $L_{0}$, are linear combinations of $\lambda^{(n_{f})}=\alpha_{\text{S}}\beta^{(n_{f})}_{0}\log{\frac{1}{v}}$, $\lambda^{(n_{f})}_{\text{cut}}=\alpha_{\text{S}}\beta^{(n_{f})}_{0}\log{\frac{1}{z_{\text{cut}}}}$, and $\lambda^{(n_{f})}_{\xi_{i}}$, defined above. In contrast, $c$ can only assume the values: $\left\{\frac{1}{1+\beta},-\frac{\beta}{1+\beta},-1,0\right\}$.
The integral in Eq. (A) can be written as:
@@
The results in Eq. (63) are valid for $n_{f}=5,4,3$. When $k_{t}$ becomes smaller than the non-perturbative scale $\Lambda$, we freeze the coupling at $\Lambda$, see Eq. (A). In this case the result of Eq. (62) is straightforward, and reads:
@@ Appendix B Mapping out the different regions
Clearly, we have to take into account additional constraints that will very much reduce the number of possible cases much below $7!$. We have
@@ 3 $z_{c}^{2}\xi_{b}^{1+\beta}$ $z_{c}^{2}\xi_{c}^{1+\beta}$ $\xi_{\Lambda}$
Thus, we have to consider 14 cases. This is what actually happens for light jets and $c$-jets. For $b$-jets, the situation is slighly simpler because the dead cone at $\theta_{i}^{2}=\xi_{b}$ implies that we are not sensitive to the scale $z_{c}^{2}\xi_{c}^{1+\beta}$. Thus, we have only 3 distinct subcases for (a) and (b), 2 for (c) and still 1 for (d), totalling 9 cases for $b$-jets.
In this work, we have fixed the value of $z_{c}=0.1$ and considered only two distinct values for the angular exponent $\beta=0,1$. We have also worked with just one value for the anti-$k_{t}$ jet radius, $R_{0}=0.4$, and fixed the non-perturbative scale $\Lambda=1$ GeV. This results in only two different hierarchies to be considered considered. In fact, for $p_{t}=150,300$ GeV we have:
$\begin{split}z_{\text{cut}}^{2}>\xi_{b}>\xi_{c}>\xi_{\Lambda}>z_{\text{cut}}^{2}\xi_{b}^{1+\beta}>z_{\text{cut}}^{2}\xi_{c}^{1+\beta}>z_{\text{cut}}^{2}\xi_{\Lambda}^{1+\beta},\end{split}$ (70)
@@
$\xi_{b}>z_{\text{cut}}^{2}>\xi_{c}>\xi_{\Lambda}>z_{\text{cut}}^{2}\xi_{b}^{1+\beta}>z_{\text{cut}}^{2}\xi_{c}^{1+\beta}>z_{\text{cut}}^{2}\xi_{\Lambda}^{1+\beta}.$ (71)
@@ Appendix C Results for the $\theta_{g}$ distribution at low and high transverse momentum
We collect here results for the $\theta_{g}$ distribution at NLL, for two different values of the transverse momentum cut of the jet. In Fig. 6, we show the case $p_{t}\geq 50$ GeV, while in Fig. 7 we consider a high-$p_{t}$ cut, nameley $p_{t}\geq 300$ GeV. Details of the plots are the same as in the main text, Fig. 2.

### 29 Feb 2024 v2

@@ Appendix A Running coupling integrals with quark-mass thresholds
Throughout the paper, we have made use of the strong coupling in the Catani-Marchesini-Webber (CMW) scheme Catani:1990rr :
@@
The limits of integration, as well as $L_{0}$, are linear combinations of $\lambda^{(n_{f})}=\alpha_{\text{S}}\beta^{(n_{f})}_{0}\log{\frac{1}{\vartheta_{g,i}^{2}}}$, $\lambda^{(n_{f})}_{\text{cut}}=\alpha_{\text{S}}\beta^{(n_{f})}_{0}\log{\frac{1}{z_{c}}}$, $\lambda^{(\text{NP})}=\alpha_{\text{S}}\beta_{0}^{(3)}\log{\frac{\left(p_{t}R_{0}\right)^{2}}{\Lambda^{2}}}$ and $\lambda^{(n_{f})}_{\xi_{i}}$, defined above. In contrast, $c$ can only assume the values: $\left\{\frac{1}{1+\beta},-\frac{\beta}{1+\beta},-1,0\right\}$.
As long as $k_{t}^{2}>\Lambda^{2}$, the integral in Eq. (A) can be written as:
@@
When $k_{t}$ becomes smaller than the non-perturbative scale $\Lambda$, we freeze the coupling at $\Lambda$, see Eq. (A). In this case the result of Eq. (A) is straightforward, and reads:
@@ Appendix B Mapping out the different regions
Clearly, we have to take into account additional constraints that will reduce the number of possible cases much below $7!$. We have
@@ 3 $z_{c}^{2}\xi_{b}^{1+\beta}$ $z_{c}^{2}\xi_{c}^{1+\beta}$ $\xi_{\Lambda}$
Thus, we have to consider 14 cases. This is what actually happens for light jets and $c$-jets. For $b$-jets, the situation is slightly simpler because the dead cone at $\theta_{i}^{2}=\xi_{b}$ implies that we are not sensitive to the scale $z_{c}^{2}\xi_{c}^{1+\beta}$. Thus, we have only 3 distinct subcases for (a) and (b), 2 for (c) and still 1 for (d), totalling 9 cases for $b$-jets.
In this work, we have fixed the value of $z_{c}=0.1$ and considered only two distinct values for the angular exponent $\beta=0,1$. We have also worked with just one value for the anti-$k_{t}$ jet radius, $R_{0}=0.4$, and fixed the non-perturbative scale $\Lambda=1$ GeV. This results in only two different hierarchies to be considered. In fact, for $p_{t}=150,300$ GeV we have:
$\begin{split}z_{c}^{2}>\xi_{b}>\xi_{c}>\xi_{\Lambda}>z_{c}^{2}\xi_{b}^{1+\beta}>z_{c}^{2}\xi_{c}^{1+\beta}>z_{c}^{2}\xi_{\Lambda}^{1+\beta},\end{split}$ (70)
@@
$\xi_{b}>z_{c}^{2}>\xi_{c}>\xi_{\Lambda}>z_{c}^{2}\xi_{b}^{1+\beta}>z_{c}^{2}\xi_{c}^{1+\beta}>z_{c}^{2}\xi_{\Lambda}^{1+\beta}.$ (71)
@@ Appendix C Results for the $\theta_{g}$ distribution at low and high transverse momentum
We collect here results for the $\theta_{g}$ distribution at NLL, for two different values of the transverse momentum cut of the jet. In Fig. 6, we show the case $p_{t}\geq 50$ GeV, while in Fig. 7 we consider a high-$p_{t}$ cut, namely $p_{t}\geq 300$ GeV. Details of the plots are the same as in the main text, Fig. 2.